• Find the volume of the solid formed by revolving the curve, r=a(1+cosθ), about the initial line.
1
Expert's answer
2021-06-14T15:07:49-0400
"V=\displaystyle\int_{0}^{\pi}\dfrac{2}{3}\pi r^2r\sin\theta d\theta"

"=\displaystyle\int_{0}^{\pi}\dfrac{2}{3}\pi (a(1+\cos \theta))^3\sin\theta d\theta"

"=\dfrac{2}{3}\pi a^3\big[-\dfrac{1}{4}(1+\cos \theta)^4\big]\begin{matrix}
 \pi \\
 0
\end{matrix}"

"=-\dfrac{1}{6}\pi a^3(0-2^4)"

"=\dfrac{8}{3}\pi a^3 (units^3)"

"V=\dfrac{8}{3}\pi a^3 \text{ cubic units}"